What is the value of cos245- sin245?
please answer this question with brief explination
 

Grade:12

3 Answers

Phanindra
6 years ago
a^2 - b^2 = (a+b)(a-b)....So... However you will get the answer without applying the formula...Cos45 = 1/root2 .. so.. cos^2 45 = 1/2Sin45 = 1/root2..so..sin^2 45 = 1/2 ...So....required answer is 0...Hope you have understood the answer..
Soumendu Majumdar
6 years ago
Dear Student,
cos 45 = sin 45 = 1/\sqrt[]{2}
so cos^{2} 45 = sin^{2} 45 = ½
so cos^{2} 45 –  sin^{2} 45 = ½ – ½ = 0 (Ans)
If it hadn’t been 45 degrees and some other angle for which you don’t know the value of cos theta and sin theta
you can use the formula cos^{2}\Theta - sin^{2}\Theta = cos2\Theta
 
Hope it helps!
 
 
Soumendu Majumdar
6 years ago
If you have any doubts, feel free to ask
I will be happy to help you!
And these formulas of trigonometry can be found in any maths book of class 11.
You can refer to the book to understand the formulas and it's uses
 
with regards,
Soumendu